# Parrot: Pareto-optimal Multi-Reward Reinforcement Learning Framework for Text-to-Image Generation

### v1

@@ Abstract
Recent works demonstrate that using reinforcement learning (RL) with quality rewards can enhance the quality of generated images in text-to-image (T2I) generation. However, a simple aggregation of multiple rewards may cause over-optimization in certain metrics and degradation in others, and it is challenging to manually find the optimal weights. An effective strategy to jointly optimize multiple rewards in RL for T2I generation is highly desirable. This paper introduces Parrot, a novel multi-reward RL framework for T2I generation. Through the use of the batch-wise Pareto optimal selection, Parrot automatically identifies the optimal trade-off among different rewards during the RL optimization of the T2I generation. Additionally, Parrot employs a joint optimization approach for the T2I model and the prompt expansion network, facilitating the generation of quality-aware text prompts, thus further enhancing the final image quality. To counteract the potential catastrophic forgetting of the original user prompt due to prompt expansion, we introduce original prompt centered guidance at inference time, ensuring that the generated image remains faithful to the user input. Extensive experiments and a user study demonstrate that Parrot outperforms several baseline methods across various quality criteria, including aesthetics, human preference, image sentiment, and text-image alignment.
Figure 1: Parrot visual examples. Parrot consistently improves the quality of generated images across multiple criteria: aesthetics, human preference, text-image alignment, and image sentiment. Moreover, Parrot handles prompt expansion to complement semantics of the text input for high-quality image generation. Each column generated images using the same seed.
* This work was done during an internship at Google.† This work was done during working at Google.
@@ 1 Introduction
Recent progress in diffusion models [15, 45, 55, 26, 36] combined with pre-trained text encoders has resulted in remarkable results in text-to-image (T2I) generation. These advancements have enabled the creation of high-fidelity novel images based on user-input text prompts. However, despite these advances, existing T2I models, such as Imagen [41] and Stable Diffusion [38], still generate images that fall short of the ideal. For example, the images in the first row in Fig. 1 exhibits quality issues such as poor composition (e.g. bad cropping), misalignment with the user input (e.g. missing objects), or generally less aesthetic pleasing. Consequently, achieving higher quality generated images with T2I generation remains a challenging task.
Quality of text-to-image generation is affected by multiple factors. Diverse quality metrics have been proposed to quantify each of these quality aspects. Some quality aspects are universal to images, including the aesthetics [21, 20, 31] of an image or whether the image is pleasing to look at [43]. There are also unique aspects regarding generated images, such as image-text alignment [33] and human preference [23] given the input prompt. Recent works [4, 10, 24] have demonstrated that incorporating such quality signals as a reward function can enhance the generated image quality of T2I models. In particular, DPOK [10] shows online reinforcement learning (RL) using ImageReward [51] score as a reward is able to improve T2I generated image quality. Given the aforementioned diverse criteria for the quality of generated images, it is a natural extension to incorporate multiple quality metrics in the RL fine-tuning of T2I models. While a straightforward strategy involves using a weighted sum of the reward scores [44, 30, 37, 49], this approach requires manual tuning of weights for different rewards, which is both time-consuming and resource-intensive. It is also not scalable as the number of rewards increase. Furthermore, optimizing one quality metric may inadvertently lead to degradation in others. For example, the model might generate an exceptionally good image without considering the relevance to the user provided prompt. Therefore, an effective strategy is needed to identify the optimal trade-off between different rewards in T2I generation.
To achieve this, we propose a novel Pareto-optimal multi-reward reinforcement learning framework for text-to-image generation, denoted as Parrot. Within a batch of samples produced by T2I models, each sample embodies a distinctive trade-off among various reward functions. By identifying and leveraging the set that attains the optimal trade-off (i.e. the Pareto-optimal set [29]) within such a training batch, Parrot effectively optimize multiple rewards simultaneously. This results in the generation of images with good aesthetics, proper image-text alignment, alignment with human preference, and an overall pleasing sentiment (Fig. 1).
The quality of generated images is significantly influenced by the text prompt input provided to the T2I model. Semantically rich prompts have been shown to produce higher quality images. As an example, simple prompts like “A boat” often fail to generate highly detailed images. On the other hand, prompts with details, like “A green boat floating on the water with a beautiful sunset”, tend to generate higher quality images. Recognizing the difficulty of manually crafting effective prompts, Promptist [13] has been introduced to autonomously learn prompt expansion with large language models (LLM). This involves RL-tuning of the LLMs while keeping the T2I model frozen as a black box. However, since the T2I model is not tuned in collaboration with the prompt expansion network, it may struggle to adapt to the generated text inputs. In Parrot, we propose to jointly optimize the prompt expansion network and the T2I model using multiple quality rewards. This allows the prompt expansion network and T2I model to collaboratively generate higher quality images. To the best of our knowledge, we are the first to show the potential of joint optimization for prompt expansion and T2I models.
Moreover, although prompt expansion facilitates a more detailed context, there is a risk of catastrophic forgetting of the original input prompt. To address this challenge, we further introduce original prompt-centered guidance in Parrot during inference time, enabling the generated images to remain faithful to the original prompts while simultaneously incorporating additional details.
Our contributions can be summarized as follows:
@@
  • We present Parrot, a novel framework that enhances the generated image quality of T2I generation through multi-reward RL fine-tuning. Leveraging batch-wise Pareto-optimal selection, it efficiently optimizes the T2I model with multiple rewards, enabling collaborative improvement in aesthetics, human preference, image sentiment, and text-image alignment.
  • We are the first to advocate for the joint optimization of the prompt expansion network and the T2I model. Our results demonstrate the superiority of this joint optimization compared to the individual fine-tuning of either the prompt expansion network or the T2I model.
  • We propose original prompt-centered guidance as a means to prevent the catastrophic forgetting of the initial input prompt after prompt expansion. This ensures that the generated image stays true to the original prompt while adding more details.
@@
  • Extensive results and a user study show that Parrot outperforms several baseline methods across various quality criteria.
@@ 2 Related Work
T2I Generation: The goal of T2I generation is to create an image given an input text prompt. Several T2I generative models have been proposed and have demonstrated promising results [5, 54, 35, 41, 18, 12, 25, 40, 19, 7]. Notably, Stable Diffusion [38] models show impressive generation performance in T2I generation, leveraging latent text representations from LLMs. Despite substantial progress, the images generated by those models still exhibit quality issues, such as bad cropping or misalignment with the input texts.
RL for T2I Fine-tuning: Fan et al. [9] pioneered the application of RL fine-tuning for T2I models, treating the denoising process as a multi-step decision making task. For human preference learning, recent works [4, 10, 6] have explored RL fine-tuning technique for T2I diffusion model, showcasing superior performance. In addition to fine-tuning the T2I model directly using RL, Promptist [13] fine-tunes the prompt expansion model by using image-text alignment scores and aesthetic scores as rewards. However, since the T2I model is not jointly fine-tuned, it may not adapt well to the following image generation tasks. Parrot proposes to jointly optimize the prompt expansion model and the T2I model using multi-reward RL.
@@
Generated Image Quality: The quality assessment of images generated by T2I models involves multiple dimensions, and various metrics have been proposed. In this paper, we consider using four types of quality metrics as rewards:  aesthetics, human preference, image-text alignment, and image sentiment. Aesthetics, capturing the overall visual appeal of the image, has been addressed using learning-based that leverage human ratings for aesthetics in real images  [31, 20, 21, 11, 16, 52, 48]. Human preferences, rooted the concept of learning from human feedback [3, 32], involves gathering preferences at scale by having raters to compare different generations [23, 51]. Text-image alignment measures the extent to which the generated image aligns with the input prompt, CLIP [33] score is often employed, measuring the cosine distance of between contrastive image embedding and text embedding. Image sentiment is important for ensuring the generated image evokes positive emotions in the viewer. Serra et al. [43] predict average polarity of sentiments an image elicits, learning estimates for positive, neutral, and negative scores. In Parrot, we use its positive score as a reward for positive emotions.
@@ 3 Preliminary
where the pre-trained diffusion model $p_{\theta}$ produces a sample distribution $p_{\theta}({\bf x}_{0}|c)$ using text condition $c$. Modifying this equation, Lee et al. [10] demonstrate that the gradient of objective function $\nabla\mathcal{J}_{\theta}$ can be calculated through gradient ascent algorithm without using the gradient of reward model as follows:
@@
where $T$ denotes the total time step of the diffusion sampling process. With parameters $\theta$, the expectation value can be taken over trajectories of the diffusion sampling process.
@@ 4.1 Parrot Overview
Figure 2: Overview of the Parrot framework. During the training, $N$ images are generated from the T2I model using the expanded prompt from the prompt expansion network. Multiple quality rewards are calculated for each image, and the Pareto-optimal set is identified using the non-dominated sorting algorithm. These optimal images are then used to perform policy gradient update of the parameters of T2I diffusion and prompt expansion network jointly. During the inference, both the original prompt and the expanded prompt are provided to the T2I model, enabling a trade-off between faithfulness and detail.
@@ 4.2 Batch-wise Pareto-optimal Selection
Algorithm 1 outlines the procedure of Parrot. Rather than updating the gradients using all images, Parrot focuses on high-quality samples, considering multiple quality rewards in each mini-batch. In the multi-reward RL, each sample generated by the T2I model presents distinct trade-offs for each reward. Among these samples, a subset with varied optimal trade-offs across objectives, known as the Pareto set, exists. For a Pareto-optimal sample, none of its objective values can be further improved without damaging others. In other words, the Pareto-optimal set is not dominated by any data points, also known as the non-dominated set. To achieve a Pareto-optimal solution with T2I diffusion model, Parrot selectively uses data points from the non-dominated set using non-dominated sorting algorithm. This naturally encourages the T2I model to produce Pareto-optimal samples with respect to the multi-reward objectives.
Reward-specific Preference: Inspired by the use of preference information in multi-objective optimization [27], Parrot incorporates the preference information through reward-specific identifies. This enables Parrot to automatically determine the importance for each reward objective. Concretely, we enrich the expanded prompt $\hat{c}$ by prepending reward-specific identifier  $<$reward k$>$ for $k$-th reward. Based on this reward-specific prompt, $N$ images are generated and are used for maximizing the corresponding $k$-th reward model during gradient update. At inference time, a concatenation of all the reward identifiers  $<$reward 1$>$,…,$<$reward K$>$” is used for image generation.
@@
Input: Text prompt $c$, Batch size $N$, Total iteration $E$, the number of rewards: $K$, Prompt expansion network $p_{\phi}$, T2I diffusion model: $p_{\theta}$, Total diffusion time step $T$, Non-dominated set: $\mathcal{P}$
@@
         $\mathcal{P}\leftarrow$ NonDominatedSet($\{{\bf x}_{0}^{1},...,{\bf x}_{0}^{N}\}$)
         for $j$ in $\mathcal{P}$ do
             $\nabla\mathcal{J}_{\theta}\mathrel{+}=\sum_{t=1}^{T}r_{k}({\bf x}_{0}^{j},\hat{c})\cdot\nabla\log p_{\theta}(x^{j}_{t-1}|c_{k},t,x^{j}_{t})$
@@
Output: Fine-tuned diffusion model $p_{\theta}$,
            prompt expansion network $p_{\phi}$
function NonDominatedSet($\{{\bf x}_{0}^{1},...,{\bf x}_{0}^{N}\}$)
@@
Non-dominated Sorting: Parrot constructs Pareto set with non-dominated points based on trade-offs among multiple rewards. These non-dominated points are superior to the remaining solutions and are not dominated by each other. Formally, the dominance relationship is defined as follows: the image ${\bf x}_{0}^{a}$ dominates the image ${\bf x}_{0}^{b}$, denoted as ${\bf x}_{0}^{b}<{\bf x}_{0}^{a}$, if and only if $R_{i}({\bf x}_{0}^{b})\leq R_{i}({\bf x}_{0}^{a})$ for all $i\in{1,...,m}$, and there exists $j\in{1,...,m}$ such that $R_{j}({\bf x}_{0}^{a})<R_{j}({\bf x}_{0}^{b})$. For example, given the $i$-th generated image ${\bf x}_{0}^{i}$ in a mini-batch, when no point in the mini-batch dominates ${\bf x}_{0}^{i}$, it is referred to as a non-dominated point.
@@
$\nabla\mathcal{J}_{\theta}={1\over n(\mathcal{P})}\sum_{i=1,{\bf x}_{0}^{i}\in\mathcal{P}}^{N}\sum_{k=1}^{K}\sum^{T}_{t=1}r_{k}({\bf x}_{0}^{i},c_{k})\times$ (4)
$\nabla_{\theta}\log p_{\theta}({\bf x}_{t-1}^{i}|c_{k},t,{\bf x}_{t}^{i}),$
@@ 4.3 Original Prompt Centered Guidance
$\bar{\epsilon}_{\theta}=w_{1}\cdot\epsilon_{\theta}({\bf x}_{t},t,c)+(1-w_{1})\cdot\epsilon_{\theta}({\bf x}_{t},t,\text{null})+$ (5)
$w_{2}\cdot\epsilon_{\theta}({\bf x}_{t},t,\hat{c})+(1-w_{2})\cdot\epsilon_{\theta}({\bf x}_{t},t,\text{null}),$ (6)

### v2

@@ Abstract
Recent works have demonstrated that using reinforcement learning (RL) with multiple quality rewards can improve the quality of generated images in text-to-image (T2I) generation. However, manually adjusting reward weights poses challenges and may cause over-optimization in certain metrics. To solve this, we propose Parrot, which addresses the issue through multi-objective optimization and introduces an effective multi-reward optimization strategy to approximate Pareto optimal. Utilizing batch-wise Pareto optimal selection, Parrot automatically identifies the optimal trade-off among different rewards. We use the novel multi-reward optimization algorithm to jointly optimize the T2I model and a prompt expansion network, resulting in significant improvement of image quality and also allow to control the trade-off of different rewards using a reward related prompt during inference. Furthermore, we introduce original prompt-centered guidance at inference time, ensuring fidelity to user input after prompt expansion. Extensive experiments and a user study validate the superiority of Parrot over several baselines across various quality criteria, including aesthetics, human preference, text-image alignment, and image sentiment.
Figure 1: Parrot visual examples. Parrot consistently improves the quality of generated images across multiple criteria: aesthetics, human preference, text-image alignment, and image sentiment. Each column shows generated images using the same seed.
* This work was done during an internship at Google.
† This work was done during working at Google.
@@ 1 Introduction
Despite significant advancements in text-to-image (T2I) generation  [17, 45, 55, 29, 38], recent work like Imagen [42] and Stable Diffusion [39], still struggle to produce high quality images. Images in the first row in Fig. 1 illustrates such quality issues in Stable Diffusion [39], including poor composition (e.g. bad cropping), misalignment with input prompts (e.g. missing objects), or overall lack of aesthetic appeal. Assessing the quality of generated images can involve various metrics such as aesthetics [24, 23, 33], human preference [26], text-image alignment [35], and emotional appeal [44]. Enhancing T2I generation across multiple quality metrics remains a challenging task.
Recent works [4, 12, 27, 15] have demonstrated that incorporating quality signals as reward functions in fine-tuning T2I with reinforcement learning (RL) can improve image quality. For instance, Promptist [15] fine-tunes prompt expansion model using RL with the sum of aesthetics and text-image alignment scores. However, the simple weighted sum approach may not effectively handle trade-offs among multiple quality metrics. As the number of rewards increases, manually adjusting reward weights becomes impractical. Moreover, optimizing one quality metric may inadvertently compromise others, as the model might prioritize aesthetics over relevance to the input prompt. Additionally, the trade-off for different reward is not controllable after training.
To address these challenges, we propose Parrot, a Pareto-optimal multi-reward reinforcement learning algorithm to improve text-to-image generation. Unlike previous approaches that treat T2I reward optimization as a single objective optimization problem, Parrot tackles this challenge through multi-objective optimization and introduces an effective multi-reward optimization strategy to achieve Pareto optimal approximation. Intuitively, each generated sample in a batch embodies a distinctive trade-off among various quality rewards, with some samples exhibiting superior trade-offs compared to others. Instead of updating gradients using all batch samples, Parrot uses non-dominated points [32] which have better trade-offs. Consequently, Parrot automatically learns from the optimal trade-off among different rewards. Moreover, Parrot learns reward-specific preference prompts, which can be utilized individually or in combination to control the trade-off among different rewards during inference time. Unlike prior work, which either solely fine-tunes the T2I model [12] or only tunes the prompt expansion network while freezing the T2I model [15], we employ the Parrot multi-reward optimization algorithm to jointly optimize both the T2I model and the prompt expansion network. This collaborative optimization unlocks the full potential of Parrot by encouraging both more details from added context from the prompt expansion model, and the overall quality improvement on the T2I generation. During inference, we further introduce original prompt-centered guidance to ensure the output image is relevant to input prompts after prompt expansion.
In summary, our contributions can be listed as follows:
@@
  • We propose Parrot, a novel multi-reward optimization algorithm for T2I RL fine-tuning. Leveraging batch-wise Pareto-optimal selection, it effectively optimizes multiple T2I rewards, enabling collaborative improvement in aesthetics, human preference, image sentiment, and text-image alignment and also allowing to control the trade-off of different rewards using a reward related prompt during inference.
  • We show the advantage of jointly optimizing both the prompt expansion network and the T2I model, which has never been explored before.
  • We introduce original prompt-centered guidance during inference time after prompt expansion, ensuring better alignment with the original prompt while enriching image details.
@@
  • Extensive results and a user study validate that Parrot outperforms several baseline methods across various quality criteria.
@@ 2 Related Work
T2I Generation: The goal of T2I generation is to create an image given an input text prompt. Several T2I generative models have been proposed and have demonstrated promising results [5, 54, 37, 42, 21, 14, 28, 41, 22, 7]. Stable Diffusion [39] shows impressive generation performance in T2I generation, leveraging latent text representations from LLMs. Despite substantial progress, the images generated by those models still exhibit quality issues, such as bad cropping or misalignment with the input texts.
RL for T2I Fine-tuning: Starting by Fan et al. [11] to explore RL fine-tuning for T2I models, following works [4, 12, 9, 16, 8] have explored RL fine-tuning technique for T2I diffusion model, showcasing superior performance for human preference learning. DPOK [12] improves quality through RL using ImageReward [51] score as a reward with a few prompts. In addition to fine-tuning the T2I model directly using RL, Promptist [15] fine-tunes the prompt expansion model by using a simple sum of aesthetic and text-image alignment scores as reward. DRaFT [6] proposed not only differentiable rewards for efficient fine-tuning but also effectiveness of using linear summation of multi-rewards. These methods treat T2I RL as a single-objective optimization problem, while Parrot employs multi-objective optimization. Additionally, prior approaches either fine-tune the T2I model or the prompt expansion model while freezing the other. In contrast, Parrot proposes joint optimization of the prompt expansion model and the T2I model using multi-reward RL to foster better collaboration.
@@
Generated Image Quality: The quality assessment of images generated by T2I models involves multiple dimensions, and various metrics have been proposed. In this paper, we consider using four types of quality metrics as rewards: aesthetics, human preference, text-image alignment, and image sentiment. Aesthetics captures the overall visual appealingness of the image, and it is learned using human ratings for aesthetics in real images  [33, 23, 24, 13, 19, 52, 48]. Human preferences, rooted the concept of learning from human feedback [3, 34, 50], involves gathering preferences at scale by having raters to compare generated images [26, 51]. Text-image alignment measures the extent to which the generated image aligns with the input prompt, CLIP [35] score is often employed, measuring the cosine distance of between contrastive image embedding and text embedding. Image sentiment is important for ensuring the generated image evokes positive emotions in the viewer. Serra et al. [44] predict average polarity of sentiments an image elicits and learn estimates for positive, neutral, and negative scores. In Parrot, we use its positive score as a reward for positive emotions.
@@ 3 Preliminary
where the pre-trained diffusion model $p_{\theta}$ produces a sample distribution $p_{\theta}({\bf x}_{0}|c)$ using text condition $c$. Modifying this equation, Fan et al. [12] demonstrate that the gradient of objective function $\nabla\mathcal{J}_{\theta}$ can be calculated through gradient ascent algorithm without using the gradient of reward model as follows:
@@
where $T$ denotes the total time step of the diffusion sampling. With parameters $\theta$, the expectation value can be taken over the trajectories of diffusion sampling.
@@ 4.1 Parrot Overview
Figure 2: Overview of Parrot. During the training, $N$ images are sampled from the T2I model using the expanded prompt from the prompt expansion network. Multiple quality rewards are calculated for each image, and the Pareto-optimal set is identified using the non-dominated sorting algorithm. These optimal images are then used to perform policy gradient update of the parameters of T2I model and prompt expansion network jointly. During the inference, both the original prompt and the expanded prompt are provided to the T2I model, enabling better faithfulness while adding detail.
@@ 4.2 Batch-wise Pareto-optimal Selection
Li et al. [30] has demonstrated that using batchwise Pareto-set learning and selecting good samples in a batch can approximate Pareto-optimality across multiple objectives. Backed up by this theory, we propose to select non-dominated points in a batch for policy gradient update in RL to achieve Pareto-optimality. Algorithm 1 outlines the procedure of Parrot. Rather than updating the gradients using all images, Parrot focuses on high-quality samples, considering multiple quality rewards in each mini-batch. In the multi-reward RL, each sample generated by the T2I model presents distinct trade-offs for different rewards. Among these samples, a subset with varied optimal trade-offs across multiple objectives, also known as the Pareto set, exists. For a Pareto-optimal sample, none of its objective values can be further improved without damaging others. In other words, the Pareto-optimal set is not dominated by any data points, also known as the non-dominated set. To achieve a Pareto-optimal solution with text-to-image generation diffusion model, Parrot selectively uses data points from the non-dominated set using non-dominated sorting algorithm. This naturally encourages the T2I model to produce Pareto-optimal samples with respect to the multi-reward objectives.
Reward-specific Preference: Inspired by the use of preference information in multi-objective optimization [30], Parrot incorporates the preference information through reward-specific identifiers. This enables Parrot to automatically determine the importance for each reward objective. Concretely, we enrich the expanded prompt $\hat{c}$ by prepending reward-specific identifier  $<$reward k$>$ for $k$-th reward. Based on this reward-specific prompt, $N$ images are generated and are used for maximizing the corresponding $k$-th reward model during gradient update. At inference time, a concatenation of all the reward identifiers  $<$reward 1$>$,…,$<$reward K$>$” is used for image generation.
@@
Algorithm Parrot: Pareto-optimal Multi-Reward RL
Input: Prompt $c$, Batch size $N$, Total iteration $E$, the number of rewards: $K$, Prompt expansion network $p_{\phi}$, T2I diffusion model: $p_{\theta}$, Total diffusion time step $T$, Non-dominated set: $\mathcal{P}$
@@
       $\mathcal{P}\leftarrow$ NDSet($\{{\bf x}_{0}^{1},...,{\bf x}_{0}^{N}\}$)
@@
function NDSet($\{{\bf x}_{0}^{1},...,{\bf x}_{0}^{N}\}$)
@@
Output: Fine-tuned diffusion model $p_{\theta}$, prompt expansion network $p_{\phi}$
Non-dominated Sorting: Parrot constructs Pareto set with non-dominated points based on trade-offs among multiple rewards. These non-dominated points are superior to the remaining solutions and are not dominated by each other. Formally, the dominance relationship is defined as follows: the image ${\bf x}_{0}^{a}$ dominates the image ${\bf x}_{0}^{b}$, denoted as ${\bf x}_{0}^{b}<{\bf x}_{0}^{a}$, if and only if $R_{i}({\bf x}_{0}^{b})\leq R_{i}({\bf x}_{0}^{a})$ for all $i\in{1,2,...,m}$, and there exists $j\in{1,2,...,m}$ such that $R_{j}({\bf x}_{0}^{b})<R_{j}({\bf x}_{0}^{a})$. For example, given the $i$-th generated image ${\bf x}_{0}^{i}$ in a mini-batch, when no point in the mini-batch dominates ${\bf x}_{0}^{i}$, it is referred to as a non-dominated point.
@@
$\nabla\mathcal{J}_{\theta}=\sum_{k=1}^{K}{\frac{1}{n(\mathcal{P})}}\sum_{i=1,{\bf x}_{0}^{i}\in\mathcal{P}}^{N}\sum^{T}_{t=1}r_{k}({\bf x}_{0}^{i},c_{k})\times\nabla_{\theta}\log p_{\theta}({\bf x}_{t-1}^{i}|c_{k},t,{\bf x}_{t}^{i}),$ (4)
@@ 4.3 Original Prompt Centered Guidance
$\bar{\epsilon}_{\theta}=w_{1}\cdot\epsilon_{\theta}({\bf x}_{t},t,c)+(1-w_{1}-w_{2})\cdot\epsilon_{\theta}({\bf x}_{t},t,\text{null})+w_{2}\cdot\epsilon_{\theta}({\bf x}_{t},t,\hat{c}),$ (5)